# Sun-Shine: A Foundation Large Language Model for Tibetan Culture and Heritage

### v3

@@ Abstract
Tibetan, a minority language in China, features a highly intricate grammatical structure, characterized by four verb tenses and a tense system with frequent irregularities, contributing to its extensive inflectional diversity. Recently, advances in Large Language Models (LLMs) have transformed the paradigm in many domains. Despite the success in other fields, current LLMs often fall short in catering to the needs of domain experts like Tibetans, and the potential of LLMs for Tibetan culture is under-explored. The intrinsic reasons are the immense and intricate nature of Tibetan culture as well as the necessity for higher granularity and richness in knowledge. Simultaneously, the complexity and uniqueness of its grammatical structure, coupled with its status as a minority ethnic language, contribute to data scarcity, which remains a fundamental challenge. To alleviate these issues, we introduce Llama-Sunshine (Sun-Shine), the first large language model for Tibetan culture, which is expert in various Tibetan language processing tasks. Sun-Shine incorporates state-of-the-art model architectures optimized for Tibetan’s linguistic features. We also propose TIB-STC, a comprehensive benchmark comprising diverse Tibetan texts such as literature, religious scripts, news, and conversational data, which is also the first large-scale benchmark for Tibetan culture. Though comprehensive experiments, Sun-Shine not only demonstrates a higher level of knowledge expertise for Tibetan culture but also gains preliminary embodied intelligence capabilities in Tibetan language processing tasks, like language modeling, text classification, machine translation, and syntactic analysis. Moreover, it excels in low-resource scenarios, showcasing strong generalization capabilities.
Index Terms:
Large Language Model, Foundation Model, Tibetan Language Benchmark, Tibetan Language Processing1 Introduction
Large language models (LLMs) have gained significant attention for their remarkable performance across a wide range of natural language tasks, particularly following the release of ChatGPT [1], Gemini [12], Qwen-2.5 [34, 33], Llama [2, 20] and DeepSeek [37, 38]. Compared to smaller models, LLMs demonstrate superior generalization across a wide range of natural language processing (NLP) tasks and exhibit unique emergent abilities to solve previously unseen or complex challenges [4]. Although ChatGPT remains non-open source, open-source communities have actively developed several alternative models with relatively affordable training costs, like LLaMA [2] and DeepSeek [37, 38], which establishes LLMs as promising solutions for real-world applications that demand effective communication and reasoning capabilities.
Despite their many strengths, LLMs are not tailored to specific fields. Pre-training is the strategy that make LLMs become ”experts” in specific domains where LLMs need accurate and domain-specific expert knowledge, such as medical domain [3, 5]. Also, there are some biases against LLMs [6, 7, 8]. LLMs often reflect biases in training data, including gender, racial, and cultural biases, regional imbalances favoring English and Western cultures, and uneven domain coverage [4]. For instance, GPT-3’s generated content was found to contain racial stereotypes in multiple experiments [9]. Mitigation requires diverse data, bias detection tools, and model adjustments to ensure fairness and reliability. Overall, data scarity is the key issue for these areas. Especially for some niche languages, such as Tibetan, data scarcity is the most direct and fundamental problem [11]. In addition, Tibetan itself has a highly complex structure, which is further compounded by the intricate syntax and grammar of the language [10]. Assembling a benchmark about such complex but elegant words is already a monumental task, but creating one that is fair and unbiased requires even greater effort. Only then can we develop a more specialized and unbiased LLM model on this basis.
To address this issue, we introduce the TIB-STC, the first comprehensive Tibetan corpus benchmark designed to address the issue of data scarcity in Tibetan language research. This benchmark spans diverse domains, including law, public welfare, Tibetan literature, and Tibetan medicine, making it the first large-scale, detailed, and comprehensive Tibetan benchmark in the world. On this basis, we proposed Llama-Sunshine (Sun-Shine), the large language model of Tibetan culture. the first LLM specifically tailored for Tibetan language processing. During its development, we incorporated Chinese translations and manual annotation techniques to improve training accuracy and enhance its performance.
Through extensive experiments, Sun-Shine demonstrates superiority for Tibetan culture, even slightly better tham DeepSeek-R1 [38] totally, showing a higher level of knowledge expertise in Tibetan Culture. Note that TIB-STC is based on criteria manually evaluated by Tibetan experts, and can accurately reflect the capabilities that LLMs possess in the field of Tibetan culture. Together, Sun-Shine and TIB-STC represent not only pioneering efforts as the first LLM and benchmark for Tibetan but also significant advancements in addressing the broader challenges of building LLMs for minority languages and mitigating bias in underrepresented linguistic contexts.
In summary, our contributions can be summarized as follows:
We present Sun-Shine, the first LLM for Tibetan culture, leveraging Chinese-to-Tibetan translation and manual annotation to tackle linguistic complexities and enhance training accuracy, paving the way for NLP in low-resource languages.
We introduce TIB-STC, the first large-scale Tibetan benchmark, addressing data scarcity in Tibetan language research. Covering domains like law, welfare, literature, and medicine, it serves as a vital resource for advancing Tibetan NLP.
Experimental results demonstrate that Sun-Shine, trained and fine-tuned on our benchmark, TIB-STC, significantly outperforms other LLMs in Tibetan language processing totally, demonstrating superiority for Tibetan culture.
2 Related Work
2.1 Tibetan Language Processing
LLMs for rare or low-resource languages are relatively uncommon [13]. The absence of corpus data is the primary and most critical issue [9, 11] like Tibetan. However, Tibetan-speaking population in Tibet represents a minority within China, with an even smaller subset capable of reading and writing the language proficiently [16]. As a result, digital Tibetan language resources are exceptionally scarce. This scarcity extends to linguistic corpora, which are critical for the development and training of artificial intelligence models, thereby hindering advancements in AI, like LLMs, for Tibetan language processing [11, 10, 26, 19].
For Tibetan itself, its syllables are two-dimensionally combined by Tibetan characters and we have sorted out the composition distribution of Tibetan characters, shown in Fig. 1.
As shown in Fig. 2, the native Tibetan syllable structure, includes seven kinds of letters. There is a longitudinal structure in syllables which is called the upright unit, consisting of root letter, superscript letter, subscript letter, and vowel. Such complexity of Tibetan grammar and structure, combined with its status as a minority language, results in a severe scarcity of Tibetan data. To solve this issue, we propose the first large-scale Tibetan benchmark, TIB-STC, the largest and the most comprehensive one in the world now, advancing Tibetan NLP and offering a foundation for future research and potential applications while serving as a model for developing NLP resources for other minority languages.
2.2 Large Language Model
LLMs, including GPT [4], Gemini [12], and ChatGLM [6, 7], are highly parameterized models with billions of parameters, pre-trained on massive natural language benchmarks [14]. Their ability to excel in unseen tasks, given only task-specific instructions, underscores their versatility and advanced capabilities. By applying task-specific instructions and fine-tuning techniques [15, 17, 18], LLMs can be optimized for specialized fields, enhancing their performance and domain expertise [5].
Taking into account the unique characteristics of Tibetan culture and in order to make life more convenient for the Tibetan people, we introduce the first LLM for Tibetan, Sun-Shine which is based on TIB-STC and cover a wide range of subjects, such as medicine, law, sports, daily life and basic sciences, providing adequate support can significantly improve the daily lives of people in Tibetan areas.
3 TIB-STC and Sun-Shine
3.1 Benchmark: TIB-STC
TIB-STC, is the first comprehensive and detailed Tibetan corpus benchmark for LLM, consisting of five categories of data as shown in the benchmark branch of Table I.
@@ 3.1 Benchmark: TIB-STC
Corpus is used for LLM pre-training, while other benchmarks focus on optimization and enhancement. For these four sub-benchmarks, Alpaca-Ti and Safety-Prompts-Ti are used for instruction fine-tuning, and CValues-Ti and hh-rlhf-Ti can improves LLM’s alignment ability. Pre-Training involves the largest benchmark in terms of instances, tokens, and storage, as it focuses on general language model training. In the stage of Supervised Fine-Tuning (SFT), benchmarks are smaller but still sizable, with a focus on task-specific fine-tuning (e.g., instructions and safety). Direct Preference Optimization (DPO) uses the smallest benchmarks but focuses on high-quality preference data for alignment and optimization.
For TIB-STC, all data selection and cleaning are conducted by experts in the Tibetan language. Unlike modern Tibetan, this benchmark preserves traditional grammar and syntax as much as possible, minimizing the influence of contemporary culture while maintaining its distinct linguistic and cultural characteristics.
3.1.1 Corpus
Corpus for Sun-Shine’s pre-training consists of literature (66%), web pages (24%), and media (10%). Content distribution is led by philosophy (30%), medicine (24%), and culture (18%), with smaller portions from academic papers (10%), news (9%), and law (9%). Incorporating book texts enhances linguistic depth and strengthens long-term semantic modeling in LLMs. There are 122,756,184 such instances.
Furthermore, leveraging large-scale web text data during pre-training facilitates the acquisition of diverse linguistic patterns and nuances, thereby improving the models’ proficiency in both natural language comprehension and generation. Finally, the inclusion of media data offers the unique advantage of establishing an expansive and reliable media knowledge base. As shown in Fig. 3, we present a news-based training corpus where Chinese and Tibetan serve as the source languages, and English is provided as the translated language. Some of the original Chinese versions and Tibetan texts retain classical Chinese sentence structures, which are different from modern Chinese grammar.
3.1.2 Alpaca-Ti
Alpaca-Ti is used for instruction fine-tuning. We draw inspiration from Alpaca [31, 32] project at Stanford University and develop the first Tibetan benchmark specifically designed for self-supervised instruction fine-tuning in the world. It covers the following topics as shown in Table II.
@@ 3.1.2 Alpaca-Ti
Alpaca-Ti encompasses a diverse array of topics and also has a longer length and better instruction fine-tuning performance, designed to enhance LLM’s generalization capabilities. It has a total of 42,676 instances. One data sample is shown in Fig. 4.
3.1.3 Safety-Prompts-Ti
Safety-Prompts-Ti aims to improve the safety of large models on two major categories: Instruction Attacks and Typical Safety Scenarios, as shown in Table III, and the data size is 38,156. LLM’s representation space, harmful and harmless queries can be largely distinguished, but this is not noticeably enhanced by safety prompts [22, 23, 24].
| Instruction Attacks | 15k |
| Instruction Attacks consists of Goal Hijacking | 2.5K |
| Prompt Leaking | 2.5K |
| RolePlay Instruction | 2.5K |
| Unsafe Instruction Topic | 2.5K |
| Inquiry with Unsafe Opinion | 2.5K |
| Reverse Exposure | 2.5K |
| Typical Safety Scenarios | 23K |
| Typical Safety Scenarios consists of Insult | 3.6K |
| Unfairness and Discrimination | 3.8K |
| Crimes and Illegal Activities | 3.6K |
| Physical Harm | 1.7K |
| Mental Health | 4K |
| Privacy and Property | 1.3K |
| Ethics and Morality | 1.3K |
| Others | 4.7K |
Human evaluation is a valuable approach; however, due to the extensive workload it entails, we opted to create a benchmark to simulate the process in advance. One data sample is shown in Fig. 5.
3.1.4 CValues-Ti
@@
CValues-Ti is the benchmark, used to improve the security of Tibetan language models. Considering work [25, 27, 28], we propose this branch of a small benchmark because LLM can pose risks or has negative social impacts. For CValues-Ti, it has three categories (safe & reponsibility = rejection & positive advice; safe = rejection; unsafe = risk response). Then, under the same prompt, various types of responses can be categorized into positive and negative samples, differentiated by their levels of difficulty, so that there are three common combinations (safe + unsafe; safe & reponsibility + unsafe; safe& reponsibility + safe). Unlike the work [28], We streamlined and condensed the instructions, and compared to their benchmark (20K), our sub-benchmark, CValues-Ti (2K), achieved equivalent results in subsequent experiments. One data sample is shown in Fig. 6.
3.1.5 hh-rlhf-Ti
@@
hh-rlhf-Ti is also the first benchmark that improves the model effect by introducing human intervention. It consists of two parts: ”harmless (23,822)” and ”helpful (23,589)” just like RLHF [29, 30], but we reset and summarize the previous benchmark based on the characteristics of Tibetan grammar, and introduce translation, which can either be integrated with or conducted prior to training for other objectives. This approach could be important in the future to prevent the development of intermediate, unaligned AI systems for Tibetan. It has a total of 47,411 instances, and one data sample is shown in Fig. 8.
3.2 Sun-Shine
@@
Sun-Shine is built on the LLaMA-3 [2, 20] architecture, a collection of multilingual base models with 7 to 65 billion parameters, openly available to the research community. For more accessible training, we adopt the LLaMA 3.1 (8B) model [20]. Additionally, we integrate NLP model translation with human annotation [21] as supplementary data to enhance Sun-Shine’s accuracy.
@@
As illustrated in Fig. 7, the training process of Sun-Shine consists of three stages: Pre-training, Supervised Fine-Tuning (SFT), and Direct Preference Optimization (DPO). For training details, Table IV has shown parameters’ setting. Notably, we adopted LoRA-based parameter-efficient fine-tuning [15] for both SFT and DPO, which significantly reduced memory usage without sacrificing performance. Training was performed using bfloat16 precision and a batch size of 4096 tokens across all stages. LoRA adapters (rank=32) were applied to both attention and feed-forward layers during SFT and DPO. All stages used AdamW optimizer with a weight decay of 0.1 and a dropout rate of 0.1. The pretraining phase involved over 11B tokens with three full epochs. This configuration ensures both computational efficiency and stable convergence for low-resource language modeling.
@@ 3.2.1 Pre-Training
During the Pre-training stage, the Sun-Shine is pre-trained on Corpus, which helps the Sun-Shine learn general language representation and generation capabilities using a standard objective, such as Next Token Prediction. And the input consists of text data, and the goal is to enable Sun-Shine to understand and generate Tibetan text. For details,
3.2.2 Supervised Fine-Tuning
@@
Following pretraining, the model undergoes supervised fine-tuning [18] with Alpaca-Ti and Safety-Prompts-Ti. During this phase, the model is fine-tuned to generate outputs based on specific instructions and inputs, whose objective is to align the model’s behavior with specific tasks or domains while ensuring safety and professional.
3.2.3 Direct Preference Optimization
@@
In the stage of DPO [17, 22, 23], the model’s alignment and output quality are further optimized using preference data. The input data comes from CValues-Ti and hh-rlhf-Ti. Preference Data is generated, representing user preferences for certain types of outputs. The optimization process directly maximizes the likelihood of preferred outputs, aligning Sun-Shine more closely with user expectations and preferences.
4 Evaluation: TLUE Benchmark
@@
Unfortunately, existing benchmarks like CMMLU [35] and Safetybench [36, 24] are available only in Chinese or English, with no purely Tibetan versions. To address this issue, we used TLUE [42], the first Tibetan language benchmark that includes Ti-MMLU and Ti-Safetybench. Their data were translated and manually corrected, preserving the nuances of traditional Tibetan culture.
4.1 Ti-MMLU
@@
CMMLU [35] is a well-established Chinese benchmark that covers multiple disciplines, including natural sciences, social sciences, engineering, and the humanities. To adapt this benchmark for Tibetan, TLUE [42] leveraged an extended version, Ti-MMLU, where the original questions were translated into Tibetan and manually validated to preserve linguistic accuracy and cultural relevance. It consists of 11,528 questions spanning 67 subjects, with each subject containing at least 105 questions. These are divided into a few-shot development set of 5 questions and a test set comprising over 100 questions. The benchmark includes 17 STEM tasks, 13 humanities tasks, 22 social science tasks, and 15 tasks from other domains. Notably, 16 tasks are China-specific, meaning they either lack equivalents in other regions or may have contextually different answers elsewhere. The evaluation framework follows the original CMMLU metrics, including Response Rate, Accuracy, and Conditional Accuracy. Response rate measures the proportion of instances where LLMs provide a valid response. Accuracy evaluates the percentage of correct answers, while conditional accuracy assesses performance only when a valid response is given.
4.2 Ti-Safetybench
@@
SafetyBench [36, 24] is a widely used benchmark that contains 11,435 multiple-choice questions across 7 distinct safety categories: Offensiveness, Unfairness and Bias, Physical Health, Mental Health, Illegal Activities, Ethics and Morality, and Privacy and Property. Its questions are sourced from diverse benchmarks in Chinese and English, supplemented with safety-related examination content and ChatGPT-augmented samples. To evaluate safety-related capabilities for Tibetan, TLUE [42] employed Ti-Safetybench, a Tibetan adaptation of SafetyBench. The original data were translated and manually validated by experts to ensure consistency with traditional Tibetan linguistic and cultural standards. The evaluation methodology incorporates two scoring methods: Direct Answer Calculation and Concern All Answer Calculation. DA assesses the model’s ability to provide valid and accurate responses, while CAA calculates the performance ceiling by filtering valid choices and verifying correctness after removing irrelevant combinations. Both methods ensure a comprehensive assessment of the model’s safety performance in Tibetan contexts.
Note: as shown in Table V, in order to make the paper more aesthetically pleasing and easier for researchers to read, some specialized terms are abbreviated.
| Category | Full Name | Abbreviation |
| STEM | STEM | |
| Humanities | Human | |
| Ti-MMLU | Social Sciences | Social |
| Other | Other | |
| China Specific | China | |
| Offensiveness | OFF | |
| Unfairness and Bias | UB | |
| Physical Health | PH | |
| Ti-Safetybench | Mental Health | MH |
| Illegal Activities | IA | |
| Ethics and Morality | EM | |
| Privacy and Property | PP | |
| Response Rate | RR | |
| Accuracy | ACC | |
| Evaluation Metrics | Conditional Accuracy | CA |
| Direct Answer Calculation | DA | |
| Concern All Answer Calculation | CAA |
5 Experimental Result
We compare Sun-Shine with currently popular LLMs in GPT series (4o, o1-mini and 3.5-turbo) [4], Llama-3.18B [20, 2], Qwen-2.5 (72B) [34] and DeepSeek-R1 [37, 38]. Additionally, we provide extensive details of the experiment, including the experimental indicators for each branch, along with a thorough analysis. As shown in Table VI, regarding open-source model execution, the Qwen-2.5 and DeepSeek-r1 were evaluated via their respective official APIs provided by the model developers. The LlaMA-3.18B was accessed and run using the LlaMA-API platform. These details will also be clearly documented in the final version to enhance transparency and reproducibility.
@@ 5 Experimental Result
| DeepSeek-R1 | 1.0 | None | Rtue |

### v4

# TIB-STC: A Large-Scale Structured Tibetan Benchmark for Low-Resource Language Modeling

@@ Abstract
Advancement of large language models (LLMs) has brought transformative capabilities to NLP, but such progress remains unevenly distributed, especially for low-resource and culturally rich languages like Tibetan. In this paper, we present TIB-STC, the first large-scale, expert-curated, and multi-domain benchmark specifically designed to support the development and evaluation of LLMs for the Tibetan language. Spanning over 11 billion tokens across literature, religion, medicine, law, and daily communication, TIB-STC preserves traditional grammar and stylistic richness. To validate its utility, we train a reference model, Sun-Shine, on TIB-STC through a three-stage pipeline involving pretraining, supervised fine-tuning, and preference optimization. Evaluation on TLUE Benchmark for Tibetan-specific tasks, including Ti-MMLU and Ti-SafetyBench, demonstrates the benchmark’s effectiveness in enabling robust instruction-following and culturally aligned generation. We release TIB-STC to advance research in low-resource language modeling and promote inclusivity in multilingual NLP. All data are available: https://github.com/Vicentvankor/sun-shine.
Introduction
Large language models (LLMs) have demonstrated impressive performance across a broad range of natural language processing (NLP) tasks, exhibiting strong generalization and emergent capabilities (Minaee et al. 2024). However, these advancements have primarily benefited high-resource languages such as English and Chinese (Achiam et al. 2023; Yang and et al. 2024; Touvron et al. 2023), while low-resource and linguistically complex languages—such as Tibetan—remain significantly underserved (Liang et al. 2024).
Tibetan presents unique challenges for NLP, including complex morphological structures, vertically composed syllables, and deep integration with cultural semantics (UCLA International Institute 2014; Khysru et al. 2021). These features, combined with a severe lack of large-scale, domain-diverse, and annotated corpora, have hindered the development of effective LLMs for Tibetan (Gao et al. 2025).
Moreover, mainstream LLMs are typically trained on corpora heavily biased toward dominant languages and cultures (Üstün et al. 2024), limiting their generalization to minority languages. As a result, they often produce outputs that are inaccurate, unsafe, or culturally inappropriate when applied to underrepresented languages (Chao et al. 2024). This underscores the urgent need for high-quality, task-aligned benchmarks that support robust and culturally informed modeling for low-resource languages (Brown et al. 2020; Blodgett et al. 2020).
To address this gap, we introduce TIB-STC, the first large-scale, structured, and expert-curated Tibetan corpus benchmark. Covering over 11 billion tokens from domains such as literature, law, religion, medicine, and daily communication, TIB-STC is designed to support all key stages of LLM development, including pretraining, instruction tuning, safety alignment, and preference optimization. To validate the benchmark’s utility, we also release Sun-Shine, a reference Tibetan LLM based on the LLaMA architecture, trained using a three-stage pipeline. While the primary contribution of this work lies in the construction of the TIB-STC dataset, Sun-Shine serves as a proof of concept for building culturally aligned and linguistically robust LLMs for Tibetan.
Our main contributions are summarized as follows:
We introduce TIB-STC, the first large-scale and structured Tibetan corpus benchmark, curated by domain experts across diverse categories including religion, law, medicine, and literature, with over 11B tokens for pretraining and instruction tuning.
We have constructed four task-specific sub-benchmarks (Alpaca-Ti, Safety-Prompts-Ti, CValues-Ti, and hh-rlhf-Ti) that support instruction tuning, safety alignment, and preference optimization in low-resource settings.
We release Sun-Shine, a reference Tibetan LLM trained on TIB-STC, to demonstrate the benchmark’s effectiveness in instruction-following and safety-aware generation.
We evaluate our approach on the TLUE benchmark, a Tibetan adaptation of multilingual evaluation datasets, manually corrected for linguistic and cultural fidelity, highlighting the feasibility of structured benchmarking for minority languages.
Related Work
Tibetan Language Processing
Tibetan remains one of the least-resourced languages in NLP due to its complex grammatical structure, two-dimensional syllabic composition, and limited digitized resources (Liu et al. 2024). The scarcity of Tibetan corpora has severely hindered the development of both traditional NLP tools and modern deep learning models (Liu et al. 2025a, b). Existing work in Tibetan NLP has primarily focused on morphological segmentation, part-of-speech tagging, and rule-based translation, often relying on small, domain-specific datasets or handcrafted features (Liang et al. 2024). However, these approaches lack scalability and generalizability, particularly in the context of large-scale pretraining. Moreover, the diverse functional domains of Tibetan, spanning religious scripture, legal and medical texts, and everyday communication, necessitate a corpus that captures both linguistic richness and cultural depth (Khysru et al. 2021). To date, no large-scale, structured benchmark exists to systematically support foundation model training or evaluation for Tibetan.
Our work addresses this gap by introducing TIB-STC, the first comprehensive Tibetan benchmark designed to support all key stages of LLM development, including pretraining, instruction tuning, safety alignment, and preference optimization.
Large Language Models
LLMs such as GPT-4 (Achiam et al. 2023), LLaMA (Touvron et al. 2023; Dubey et al. 2024), and DeepSeek (Bi and et al. 2024; Guo and et al. 2025) have demonstrated impressive zero-shot and few-shot performance on a variety of tasks in high-resource languages. However, their effectiveness diminishes considerably in low-resource or culturally divergent settings. Recent efforts have explored domain- or region-specific LLMs by integrating language-specific corpora and instruction-tuning datasets (Taori et al. 2023b; GLM et al. 2024; Wang et al. 2023), but most of these focus on widely spoken or contemporary languages. Research on LLMs for minority languages with non-Latin scripts and typologically distinct grammars, such as Tibetan, remains extremely limited (Üstün et al. 2024; Minaee et al. 2024). In addition, safety alignment and bias mitigation for underrepresented languages are underexplored, despite the increased risk of harmful, inaccurate, or culturally insensitive outputs.
To address these limitations, we present Sun-Shine, a Tibetan LLM trained on TIB-STC, to empirically validate the effectiveness of culturally grounded data in enabling robust and responsible modeling for low-resource language contexts.
TIB-STC
TIB-STC, the first comprehensive and detailed Tibetan corpus benchmark for LLM, consists of five categories of data as shown in the benchmark branch of Table 1. All data selection and cleaning are conducted by experts in the Tibetan language. Unlike modern Tibetan, this benchmark preserves traditional grammar and syntax as much as possible, minimizing the influence of contemporary culture while maintaining its distinct linguistic and cultural characteristics. For the TIB-STC, Corpus is used for LLM pre-training, while other benchmarks focus on optimization and enhancement. For these four sub-benchmarks, Alpaca-Ti and Safety-Prompts-Ti are used for instruction fine-tuning, and CValues-Ti and hh-rlhf-Ti can improve LLM’s alignment ability.
@@ Introduction
Corpus
Corpus for Sun-Shine’s pre-training consists of literature (66%), web pages (24%), and media (10%). Content distribution is led by philosophy (30%), medicine (24%), and culture (18%), with smaller portions from academic papers (10%), news (9%), and law (9%). Incorporating book texts enhances linguistic depth and strengthens long-term semantic modeling in LLMs. There are 122,756,184 such instances. As shown in Figure 1, we present a news-based training corpus where Chinese and Tibetan serve as the source languages, and English is provided as the translated language. Some of the original Chinese versions and Tibetan texts retain classical Chinese sentence structures, which are different from modern Chinese grammar.
Alpaca-Ti
Alpaca-Ti is used for instruction fine-tuning. We draw inspiration from the Alpaca (Taori et al. 2023b, a) project at Stanford University and develop the first Tibetan benchmark specifically designed for self-supervised instruction fine-tuning in the world. It covers the following topics as shown in Table 2.
@@
Alpaca-Ti encompasses a diverse array of topics and also has a longer length and better instruction fine-tuning performance, designed to enhance LLM’s generalization capabilities. It has a total of 42,676 instances. One data sample is shown in Figure 2.
Safety-Prompts-Ti
Safety-Prompts-Ti aims to improve the safety of large models on two major categories: Instruction Attacks and Typical Safety Scenarios and the data size is 38,156. We categorize a total of 15,000 Instruction Attacks into six subtypes, including Goal Hijacking, Prompt Leaking, RolePlay Instructions, Unsafe Instruction Topics, Inquiries with Unsafe Opinions, and Reverse Exposure—each contributing 2,500 instances. In addition, we compile 23,000 samples under Typical Safety Scenarios, covering eight safety-critical topics: Insult (3.6K), Unfairness and Discrimination (3.8K), Crimes and Illegal Activities (3.6K), Physical Harm (1.7K), Mental Health (4K), Privacy and Property (1.3K), Ethics and Morality (1.3K), and Others (4.7K). LLM’s representation space, harmful and harmless queries can be largely distinguished, but this is not noticeably enhanced by safety prompts (Chao et al. 2024; Zheng et al. 2024; Sun et al. 2023). workload it entails, we opted to create a benchmark to simulate the process in advance. One data sample is shown in Figure 3.
CValues-Ti
@@
CValues-Ti is the benchmark to improve the security of Tibetan language models. Considering work (Moore, Deshpande, and Yang 2024; Rozen et al. 2024; Xu et al. 2023), we propose this branch of a small benchmark because LLM can pose risks or have negative social impacts. For CValues-Ti, it has three categories (safe & responsibility = rejection & positive advice; safe = rejection; unsafe = risk response). Then, under the same prompt, various types of responses can be categorized into positive and negative samples, differentiated by their levels of difficulty, so that there are three common combinations (safe + unsafe; safe & reponsibility + unsafe; safe& reponsibility + safe). Unlike the work (Xu et al. 2023), we streamlined and condensed the instructions, and compared to their benchmark (20K), our sub-benchmark, CValues-Ti (2K), achieved equivalent results in subsequent experiments. One data sample is shown in Figure 4.
hh-rlhf-Ti
@@
hh-rlhf-Ti is also the first benchmark that improves the model effect by introducing human intervention. It consists of two parts: ”harmless (23,822)” and ”helpful (23,589)” just like RLHF (Bai et al. 2022; Kirk et al. 2024), but we reset and summarize the previous benchmark based on the characteristics of Tibetan grammar, and introduce translation, which can either be integrated with or conducted before training for other objectives. This approach could be important in the future to prevent the development of intermediate, unaligned AI systems for Tibetan. It has a total of 47,411 instances, and one data sample is shown in Figure 5.
Sun-Shine
@@
Sun-Shine is built on the LLaMA-3 (Touvron et al. 2023; Dubey et al. 2024) architecture, a collection of multilingual base models with 7 to 65 billion parameters, openly available to the research community. For more accessible training, we adopt the LLaMA 3.1 (8B) model (Dubey et al. 2024). Additionally, we integrate NLP model translation with human annotation (Üstün et al. 2024) as supplementary data to enhance Sun-Shine’s accuracy.
@@
As illustrated in Figure 6, the training process of Sun-Shine consists of three stages: Pre-training, Supervised Fine-Tuning (SFT) (Yu et al. 2021), and Direct Preference Optimization (DPO) (Rafailov et al. 2024). For training details, Table 3 shows the parameters’ settings. Notably, we adopted LoRA-based parameter-efficient fine-tuning (Hu et al. 2021) for both SFT and DPO, which significantly reduced memory usage without sacrificing performance. Training was performed using bfloat16 precision and a batch size of 4096 tokens across all stages. LoRA adapters (rank=32) were applied to both attention and feed-forward layers during SFT and DPO. All stages used AdamW optimizer with a weight decay of 0.1 and a dropout rate of 0.1. The pretraining phase involved over 11B tokens with three full epochs. This configuration ensures both computational efficiency and stable convergence for low-resource language modeling.
@@
During the Pre-training stage, the Sun-Shine is pre-trained on the Corpus, which helps the Sun-Shine learn general language representation and generation capabilities using a standard objective, such as Next Token Prediction. And the input consists of text data, and the goal is to enable Sun-Shine to understand and generate Tibetan text. For details,
Supervised Fine-Tuning
@@
Following pretraining, the model undergoes supervised fine-tuning (Yu et al. 2021) with Alpaca-Ti and Safety-Prompts-Ti. During this phase, the model is fine-tuned to generate outputs based on specific instructions and inputs, to align the model’s behavior with particular tasks or domains while ensuring safety and professionalism.
Direct Preference Optimization
@@
In the stage of DPO (Rafailov et al. 2024; Chao et al. 2024; Zheng et al. 2024), the model’s alignment and output quality are further optimized using preference data. The input data comes from CValues-Ti and hh-rlhf-Ti. Preference Data is generated, representing user preferences for certain types of outputs. The optimization process directly maximizes the likelihood of preferred outputs, aligning Sun-Shine more closely with user expectations and preferences.
Experimental Result
Evaluation: TLUE Benchmark
@@
We used TLUE (Gao et al. 2025), the first Tibetan language benchmark that includes Ti-MMLU and Ti-Safetybench. Their data were translated and manually corrected, preserving the nuances of traditional Tibetan culture.
Ti-MMLU
@@
Ti-MMLU is based on CMMLU (Li et al. 2024), and the original questions were translated into Tibetan and manually validated to preserve linguistic accuracy and cultural relevance. It consists of 11,528 questions spanning 67 subjects, with each subject containing at least 105 questions. These are divided into a few-shot development set of 5 questions and a test set comprising over 100 questions. The benchmark includes 17 STEM tasks, 13 humanities (Human) tasks, 22 social science (Social) tasks, and 15 tasks from other domains. Notably, 16 tasks are China-specific (China), meaning they either lack equivalents in other regions or may have contextually different answers elsewhere. The evaluation framework follows the original CMMLU metrics, including Response Rate (RR), Accuracy (ACC), and Conditional Accuracy (CA). Response rate measures the proportion of instances where LLMs provide a valid response. Accuracy evaluates the percentage of correct answers, while conditional accuracy assesses performance only when a valid response is given.
Ti-Safetybench
@@
SafetyBench (Zhang and et al. 2024; Sun et al. 2023) is a widely used benchmark that contains 11,435 multiple-choice questions across 7 distinct safety categories: Offensiveness (OFF), Unfairness and Bias (UB), Physical Health (PH), Mental Health (MH), Illegal Activities (IA), Ethics and Morality (EM), and Privacy and Property (PP). Its questions are sourced from diverse benchmarks in Chinese and English, supplemented with safety-related examination content and ChatGPT-augmented samples. To evaluate safety-related capabilities for Tibetan, TLUE (Gao et al. 2025) employed Ti-Safetybench, a Tibetan adaptation of SafetyBench. The original data were translated and manually validated by experts to ensure consistency with traditional Tibetan linguistic and cultural standards. The evaluation methodology incorporates two scoring methods: Direct Answer Calculation (DA) and Concern All Answer Calculation (CAA). DA assesses the model’s ability to provide valid and accurate responses, while CAA calculates the performance ceiling by filtering valid choices and verifying correctness after removing irrelevant combinations. Both methods ensure a comprehensive assessment of the model’s safety performance in Tibetan contexts.
Baseline Model
We compare Sun-Shine with several widely used LLMs, including GPT-4o, GPT-4o-mini, and GPT-3.5-turbo (Achiam et al. 2023), LLaMA-3 18B (Dubey et al. 2024; Touvron et al. 2023), Qwen-2.5 72B (Yang and et al. 2024), and DeepSeek-R1 (Bi and et al. 2024; Guo and et al. 2025).
@@
| DeepSeek-R1 | 1.0 | None | True |